A flywheel has a moment of inertia of 4.8 kg-m2 . What constant force is required to increase its



frequency from 3 rev/s to 6 rev/s in 7 revolutions?

1
Expert's answer
2022-03-18T09:15:33-0400
"2\pi N=\frac{\omega_f^2-\omega_i^2}{2\alpha}"

The angular acceleration

"\alpha=\frac{\omega_f^2-\omega_i^2}{4\pi N}=\pi\frac{n_f^2-n_i^2}{N}"

"\alpha=3.14\times\frac{6^2-3^2}{7}=12.1\:\rm rad/s^2"

The requrid torque

"\tau=I\alpha\\
=4.8\times 12.1=58\:\rm N\cdot m"